# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v3

@@ 5. Discussion
Our analysis shows that youth imagine future social media not as a rejection of current platforms but as a reconfiguration of their strengths. They want the visibility and reach of social networking and broadcasting platforms while also craving the intimacy and authenticity found in spaces like games, voice chat, or small-group apps. Figure 2 illustrates this blended vision: an aspirational middle ground where meaningful connection is not an afterthought but the central purpose. Building on this insight, the following subsections trace how participants moved from critiquing existing platforms toward articulating design goals, culminating in our proposal of “spatial integrity” as a framework for supporting deeper forms of online connection.
The relevance of these findings has become increasingly apparent with recent technological developments. With the introduction of technologies like Genie 3 and similar platforms, the fictional world medium of remote social connection that participants envisioned is barely fictional. The 3D, open-world, and fully customizable platforms for self-expression and organic social connection that emerged from our workshops now represent concrete possibilities for the near future, much nearer than we anticipated when we began this study. This technological trajectory underscores the timeliness of our findings and their potential to guide the development of relationship-supporting technologies that leverage these emerging capabilities (DeepMind, 2025).
5.1. From Reactionary Designs to Aspirational Design Goals
There are societal concerns about social media’s impact on youth well-being (Oth, 2023). Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023). This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality. Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023). At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024). This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connections that youth actually seek. Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
However, in our earlier work with this youth population, we noticed several recurring challenges that shaped how participants thought about social media. First, there was a strong design fixation on social media design around existing apps, both in terms of interface, such as small 2D mobile screens, endless scrolling, and uniform templates, and in terms of user experience, where visibility depends on maintaining large networks of friends and followers, often at the expense of authenticity and comfort. Second, participants showed a cognitive fixation when imagining the future of social media, struggling to see it as anything beyond its currently perceived role as a convenient but limited tool for remote social connection. Third, these fixations were compounded by generally negative experiences with social media, which led many in earlier interviews to propose reactionary fixes, like stronger privacy, the removal of infinite scrolls, or eliminating general media content such as Reels. Such cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting. The procedure was intended to serve three purposes: breaking participants free from these biases, learned helplessness, and pessimism by encouraging them to think beyond the “app” or “platform” on mobile devices; equipping them with prompts that licensed thinking beyond perceived technical and practical constraints; and creating conditions where they could explore proactive and aspirational visions of social interaction rather than simply reacting to existing frustrations. Our chosen method prioritized uncovering underlying desires and needs that young people associate with meaningful remote social interaction over producing concrete feature lists.
In line with the goals of co-design processes (Sanders and Stappers, 2008), our interest lay less in the specifics of functionality and more in uncovering what kinds of interactions and environments participants valued most. Particularly noteworthy was the fact that the workshops helped participants expand their understanding of what social platforms could be, empowering them to feel more hopeful and articulate desires they might have previously dismissed as impossible within conventional social media paradigms. This demonstrates how the design method chosen in our study offered a generative path toward hope in domains where narratives of doom are prevalent.
Our analysis contributes insights about the types of remote social interactions that youth find most meaningful and fulfilling, examining how these desired interactions align with or diverge from current platform affordances, as we detail in Section 5.2. We also identify core aspects of design that can undermine relational quality from which we develop the “spatial integrity” framework introduced in Section 5.3. Taken together, we hope this work equips designers with proactive design goals for envisioning new futures of social technologies and provides an initial framework for evaluating design integrity in relationship-supporting aspects.
5.2. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
Our findings build on prior research showing that while youth are often dissatisfied with archetypal social media platforms, their goal is not to reduce time spent connecting online. Instead, they seek spaces that support deeper, more intentional social interactions. Participants described social connections that feel genuine, fulfilling, and emotionally present and not performative. The changes in social media platform design that they sought were as follows:
  1. (1)
    Immersive Experiences Through Spatial Elements. Youth expressed a strong desire to move beyond “flat” social media toward spatially immersive experiences. They imagined platforms where relationships could unfold through presence in shared 3D environments, not just posts on a feed. They described the value of feeling physically co-present with others, such as walking together in a virtual world or talking in a space that feels alive and embodied. They also envisioned personal spaces and community-centered landscapes, such as personal homes or neighborhoods that mirror real-world social structures. Several imagined futuristic forms of presence, such as those of holographic displays or VR interfaces that make interacting feel natural and embodied, reducing the sense of separation created by screens.
  2. (2)
    Organic and Intentional Social Interactions. Participants described meaningful connections as unfolding through casual, low-pressure, and community-oriented interactions. They valued flexible and safe spaces for direct conversations, where people could gather in protected and themed rooms to get to know each other more personally. They also emphasized the importance of community gathering places and shared activity spaces, such as areas to play games, collaborate on projects, or simply hang out. Visions of these spaces encompassed both low-intensity ambient co-presence, such as sitting together quietly while listening to music, and public third places that enable serendipitous encounters with new people. Participants also described the value of effortful and intentional navigation, preferring to move through spaces purposefully rather than having content or people pushed to them by algorithms.
  3. (3)
    Expressive and Lower-Stakes Sharing. Participants sought ways to express their identities naturally and casually, without the pressure of curating formal posts. They valued personal area and avatar customization as ways to reflect their personality through how they decorate spaces or design their avatars. They also envisioned ways to share real-time states and updates, such as showing what they are doing in the moment, fostering a sense of immediacy and connection. Participants described the importance of sharing multisensory or emotional memories, imagining ways to let others experience meaningful moments, like a special memory or sensory detail, in richer and more embodied ways than existing media allows.
  4. (4)
    Granular, Intuitive, and Playful Privacy Mechanisms. Participants emphasized the need for privacy that feels natural and understandable, not overwhelming or rigid. They described space-based privacy, where the type of space, such as a public plaza or a personal bedroom, intuitively signals boundaries. They also described playful privacy mechanisms, such as secret knocks or hidden rooms, that make managing access feel engaging and enjoyable. Participants expressed the importance of contextual privacy, with the ability to adjust what different people see in different contexts or moments. They wanted to feel fully in control of their visibility, including the ability to become invisible or undiscoverable to people they block or want to avoid.
  5. (5)
    Refocusing Interaction Priorities for Meaningful Engagement. Participants expressed frustration with how mainstream platforms prioritize endless content streams, celebrity updates, and commercial content over personal connections with people they care about. They described the value of intentional content consumption, where they could choose what they wanted to engage with rather than being fed endless streams. They also advocated for reducing celebrity and commercial influence, preferring to focus on posts from close connections. Some participants described using spatial representations of relationships, such as placing their closest friends’ spaces physically nearer to them in the platform’s landscape. Across these ideas, participants emphasized the value of purposeful, meaningful use of time, where interaction feels worthwhile because it supports relationships, not just engagement metrics or passive scrolling.
These findings challenge the prevailing assumption that engaging youth requires addictive features or endless content streams. Instead, youth express a clear desire for platforms that prioritize meaningful connection over superficial engagement. While some of their specific suggestions may not be immediately implementable, their underlying desires reveal important principles for future platform design: spaces where there is a genuine presence of others, relationships can develop organically, privacy feels intuitive rather than bureaucratic, self-expression can be subtle and multifaceted, and users have genuine control over curating their social experience around relationships they truly value.
Our participants envision environments where relationships deepen organically through shared experiences and deliberate engagement rather than through metrics-driven interactions. This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement, challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
5.3. Spatial Integrity for Meaningful Social Connections in Online Environments
Concept Description Supported Interactions
Spatial Presence The ability to perceive and navigate a virtual space, including awareness of other movement, and presence. Gaming and VR platforms excel at this, whereas mainstream social media provides limited spatial affordances for awareness and intentional navigation. Immersion and embodiment (Section 4.1); memory and experience sharing (Section 4.4.3); co-existing in surreal landscapes (Section 4.6.3)
Spatial Configuration The presence of fixed/defined spatial structures that facilitate interaction. Platforms like Discord create stable communal spaces with clear channels and streams. Users value this for fostering deeper relationships. Neighborhood structures (Section 4.3.1); prioritizing closer connections (Section 4.3.3); creating quality conversational spaces (Section 4.3.4); imposing clearer boundaries (Section 4.3.5); deliberate navigation and content discovery (Section 4.5.1)
Spatial Composition The segmentation of a social network into defined areas and pathways. Clear boundaries help manage social norms and interactions, with current platforms blurring personal and public spaces. Building comfort through ambient presence (Section 4.2.2); surfacing third places (Section 4.3.2)
Spatial Depth The amount of layered, rich information and space supporting richer, more intuitive posting and commenting. 3D social media limits expression to explicit actions like posting and commenting. Leveraging spatial cues for intuitive privacy control (Section 4.1.2); indirect self-expression through customization and artifact display (Section 4.4.1); memory and experience sharing (Section 4.4.3)
Table 2. Dimensions of Spatial Integrity in sociotechnical systems in supporting meaningful social connections.
Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces:
  1. (1)
    Spatial Presence: Understanding the Space and Who’s There. Spatial presence refers to the ability to perceive and navigate within a virtual space, encompassing elements such as proximity/propinquity (knowing who is nearby), navigation (moving within the space), and awareness (understanding one’s position relative to others). While gaming platforms and social VR environments often excel at creating this sense of presence, mainstream platforms like Instagram lack basic spatial affordances, limiting users’ awareness and intentional navigation within the platform.
  2. (2)
    Spatial Composition: The Meaning of the Space. Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019). Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offer only transient content streams. Our participants emphasized the importance of dedicated areas like communal lounges or study spaces where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
  3. (3)
    Spatial Configuration: How the Space is Organized. Spatial configuration is “the segmentation of space into subunits with physical boundaries and pathways between them” (Small and Adler, 2019). Clear boundaries and pathways help people recognize territories and manage social norms. Current platforms lack such clarity; a user’s “feed” is meant to be a personal space yet often lacks clear boundaries, while profiles can be accessed without users’ awareness. This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
  4. (4)
    Spatial Depth: How Much Information Can Be Perceived. Spatial depth determines how much information users can perceive in a space. 3D spaces enable higher perceptual density, allowing users to process information more intuitively (Kumar and Benbasat, 2004; McIntire et al., 2014; Tavanti and Lind, 2001). Current platforms, limited by 2D interfaces, restrict users to explicit actions like posting and commenting rather than supporting more nuanced forms of expression. Incorporating spatial depth could enable richer self-expression and more natural interactions.
According to our participants, these dimensions serve as the foundation for designing online spaces that effectively support the kind of social connections they desire. The relationship between these dimensions and the interactions they support is laid out in Table 2.
These four dimensions of spatial integrity—presence, composition, configuration, and depth—provide a framework for designing online spaces that support meaningful social connections. By prioritizing these dimensions, platforms can create environments where the kind of meaningful interactions our participants desire can flourish. While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram. Especially as spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms. (It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).) By thoughtfully incorporating these principles, platforms can create spaces that foster genuine connection and community, aligning with the desires expressed by our participants.
5.4. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently transfer to youth who are not familiar with such environments. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
Future research should empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces. This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant. They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him, including personal details they would not share with friends. Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share. This is particularly striking given that participants were otherwise skeptical of AI. This raises interesting questions about how youth might engage with AI when it is framed within a familiar and relatable narrative. It suggests that embedding AI in natural, empathetic contexts might lead people to trust and interact with it far more than they otherwise would. Future research should delve into the privacy implications of this dynamic and explore how narrative framing could influence users’ relationships with AI in ways that are not immediately obvious.
@@ 6. Conclusion
This study explored how youth with 3D gaming experience reimagine social media when freed from current platform constraints. Through Hogwarts’ fictional world, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption defining mainstream social media.
Their visions called for platforms that feel more like 3D games than Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to real relationships’ complexity while incorporating playfulness and whimsy.
Building on these insights, we introduced spatial integrity, foundational design principles enabling such visions. While spatial integrity alone cannot guarantee meaningful connection, it provides structural conditions supporting the relational, embodied, and playful interactions participants described.
The study underscores that youth seek not less time on social media, but more fulfilling and emotionally present interactions within digital spaces. Their visions point toward a future where social media feels like the dynamic, shared social worlds found in 3D gaming environments. While our findings stem from youth familiar with 3D gaming, they offer a hopeful design direction for spaces supporting the meaningful social connections youth desire.
Appendix A FI Workshop Protocol
@@ Introduction
Appendix B Participant Interview Summary
Table 3. Key Features for Envisioned Social Media from Participant Interviews
PID Envisioned Social Media Features
P01 Personalized magical houses; Pensieve-inspired memory sharing; Interactive pet guide; Neighborhood & spatial socialization; Dynamic and context-aware avatars; Front porch for public updates, pinned messages, and casual interactions; Mailbox updates; Public areas for open interaction & Interest-based rooms for community engagement; Area-based access control; Music & mood reflection in rooms; News hub or enchanted map for real-time global updates.
P02 Pensieve-inspired memory sharing; Themed rooms for open discussions; Personal vaults; Dual-identity portal; House-elf moderator; Themed personal spaces; Ephemeral, once-view pensieve posts; Interactive room creation with spells; Vault access tracking; Serendipitous interactions via walking.
P03 Personalized magical homes; Interactive town square & community voting; Floating neighborhoods; Portable portal mirror; Pensieve-inspired memory & emotion sharing; Floating holographic interface; VR-like navigation; Selective visibility & privacy settings; Muggle guest system in selective areas with limited abilities; Shopping district & immersive commerce; Immersive media; Customizable avatars & dynamic appearance.
P04 Floating memory orbs; Personalized spaces; Enchanted books with different features; Ghostly DJ with personalized music based on moods and activities; Glowing message threads that float towards recipients; Wisp notifications; Interactive exploration in a village-like landscape.
P05 Immersive memory bubbles; Personalized spaces; Handwritten air messages; Magical jars for real-time communication; Interactive social memory town; Secret rooms for privacy; 3D commenting system; House elf companion (AI assistant).
P06 Customizable personal spaces; Co-created group spaces; Animated polaroid-style letters; Teleportation-based navigation; Memory-building activities; Interactive bulletin boards & photo albums; Privacy through magical barriers like knocking on a bookshelf or moving objects in a set sequence; House elf companion (AI assistant); Intentional visit to friends’ spaces without algorithmic scrolling; Neighborhood & spatial socialization.
P07 Holographic interface; Brain-connected communication; Immersive memory sharing; Customizable profiles; Multi-threaded conversations; Personalized content curation; Integrated multi-platform content; Custom social interaction modes.
P08 Pensieve-inspired memory sharing; Animated interactive painting; Messaging through magical letters; Color-coded wand lights for notifications; VR-like social spaces; Past interaction-based communication permissions; Customizable group interactions; Marketplace for magical item trading.
P09 Personalized rooms; Themed social spaces; Space-based control access, secret areas, & hidden objects; Magical mailbox system; Intentional content consumption; Neighborhood & spatial socialization; Customizable avatars; Customizable bulletin board for updates; Moderation via magical guardians; Room of Requirement-inspired group chats.
P10 Personalized rooms & houses; Magical mailbox; Social village with communal spaces; Enchanted messenger animals; Immersive activity view; Customizable avatars & profiles; 3D art sharing; Stores integrated with real-world services; Space & object-based access control.
P11 Personalized houses; Interactive 3D avatars; Spatial discovery & connection; Magical keys to access specific areas.
P12 Immersive 3D social spaces; AI-generated personal worlds; Collaborative project spaces; Pensieve-inspired memory sharing; Temporary, safe-sharing space for struggles and emotional reflections; Adaptive privacy & anti-scam protection; 3D interactive messaging & calls; Interest-based connection zones; Engagement-based prioritization.
P13 Personalized houses; Mutual QR-based friend verification; Dynamic room-based privacy; Shuffling access mechanism; Selective private comments; Entertainment & saved content integration; Immersive 3D movement, interaction, and environment-based socialization.
P14 Hybrid 2D-3D interface; Customizable personal rooms; Multi-layered privacy control; Teen & child protection zones; Pensieve-inspired memory sharing; Real-time guided walk through personalized content; Gaming & group spaces; Mutual unfollowing & privacy indicators; Authentic avatars.
P15 3D virtual world; Personalized homes & avatars; Teleportation; Invisibility cloaks for privacy; Magical visual effects for messages; Interactive social spaces; Streak-based engagement; Integrated marketplace & digital collectibles; House elf AI assistant.
P16 Personalized virtual homes; Party system & close-knit groups; Authentic avatars; Communal spaces for social activities; Teleportation & open-world navigation; Digital memory archiving; Dynamic virtual world customization; Space-based access control.
P17 Personalized 3D avatars; Personalized virtual homes; Interactive public spaces; Spontaneous social interactions; Integrated chat & calls; Dynamic friend spaces; Intentional social connections; Space-based access control.
P18 Teleportation into shared experiences; Customizable avatars & environments; Interactive social hubs; Personalized virtual rooms; Public virtual rooms; Holographic comments & real-time reactions; Ad integration without intrusion.
P19 Personalized homes; Virtual social activities; Selective social expansion based on interests; Immersive environment exploration; Multi-tiered social zones.
P20 Immersive, interactive holograms for music & gaming; Anonymous discussion threads; Contextual commenting system; Interest-based content; Real-time online status indicators; Personalized comment moderation.
P21 Emotion-revealing teleportation; Magic book messaging; Marauder’s map social tracking; Personalized homes; Emotion-based avatars; Selective mind reading for close friends; Compatibility-based social matching; Spontaneous conversation via walking, teleportation, and environment transformation; Immersive social events; Space-based access control.
P22 Personalized avatar; Personalized virtual houses; Public & private houses; Town square & bulletin board; Interest-based groups & activities; Integrated messaging system; Personalized profile pages; Adult & teen separation.
P23 Personalized homes; Neighborhood & social spatialization; Animated sand avatars; Communication via sand figures; Real & anonymous dual identity; Interest-based social discovery; Dynamic social spaces; Space-based access control; "Hall of Fame" memory posts.

### v4

@@ 5.1. Defamiliarization, Not Novelty: How FI Expands the Frontier of Social Media Design
5.2. Minecraft, Not Instagram: Interaction Logics Beyond the Feed
Public discourse increasingly portrays social media as fundamentally detrimental to mental health, friendship, and authentic social interaction. This narrative often equates social media with specific platforms like Instagram, TikTok, or X and the ways they happen to be designed today (OtherOtherTikTokHarmst; noauthor_2023-pn). Our study participants expressed similar perspectives, frequently describing social media as centered around superficial interactions that fail to foster meaningful connections. Yet this characterization did not match where they described actually feeling close to friends.
As illustrated in Figure 1(a), there is a disconnect between platforms that youth commonly identify as “social media” (such as Instagram and TikTok) and the platforms where they actually experience meaningful connections (such as Discord or Minecraft). The latter are not typically framed as “social media” in everyday discourse, even though they often serve social functions more directly. This boundary is not merely semantic: it shapes what becomes imaginable. If “social media” is assumed to mean a 2D feed, a profile, and an attention economy, then even well-intentioned redesign prompts tend to elicit cynicism or incrementalism. When participants reasoned from spaces they already associate with social closeness (games, chat servers, shared virtual worlds), the future of social connection expanded beyond the feed metaphor.
This misalignment between expectations and experiences is problematic because negative perceptions of social media use can diminish psychological well-being (lee2024social; kim2024privacysocialnormsystematically). Youth experience disappointment in seeking deep social connections on platforms where such interactions are scarce. As shown in Figure 1(b), social media serves diverse uses and gratifications, from information seeking and entertainment to status enhancement and network maintenance. Rather than maintaining an entirely negative view confined to certain platforms, it is crucial to understand each platform’s distinct values and limitations and to name the kinds of connection each one is organized to support.
Critiquing Instagram for not facilitating authentic, deep friendships can miss the point: while such connections might occasionally form there, it is not the platform’s central form of engagement. However, platforms are not static; their designs evolve, and user engagement patterns shift over time. Instagram, primarily used for personal branding, sometimes facilitates closer communication through private ‘finsta’ accounts. This fluidity underscores the importance of teaching (WilsonWisniewski-2012-FightingMyRegulation-m; Wisniewski-2018-PrivacyParadox-l; TheLearningNetwork-2025-WhatTeensMedia-z) youth to be more discerning about the connections they seek and what a given platform’s interaction model realistically supports, rather than reinforcing anxiety-inducing narratives that treat “social media” as a single, uniformly harmful category.
Together, these accounts clarify the contrast participants repeatedly drew between platforms they associate with “social media” and the spaces where they actually feel socially connected. When describing meaningful connection, they emphasized presence, shared activity, and low-pressure co-presence, drawing on experiences in Minecraft- and Discord-like environments rather than Instagram-style broadcast and feed consumption.
5.3. From Subjects to Authors: FI as a Translator for Youth Voice
Youth are typically positioned in social media discourse as subjects to be protected or problems to be solved: screen time to be managed, mental health to be safeguarded, misinformation vulnerability to be addressed. This framing, however well-intentioned, reproduces a power imbalance in which young people live with the consequences of platform design earlier and more intensely than any other demographic, yet rarely get to author the underlying rules. FI offers a corrective. By treating participants as theorists and inventors rather than survey respondents or focus group informants, FI shifts the locus of imagination. The result is not merely a collection of wishes, but a disciplined translation pipeline that surfaces interaction primitives (what actions exist, what counts as visibility, what is reversible, what requires consent), normative frames (what should be easy versus hard, what should be private by default, what should be communal), and actionable design directions grounded in participants’ own social experience.
This repositioning matters because asking young people to imagine within realism often reproduces the very constraints that limit design thinking. Realism already encodes corporate incentives, technical path dependencies, and fear-based narratives inherited from parents, teachers, and policymakers. When participants are prompted to “fix” social media, they inherit a frame saturated with cynicism and incrementalism. FI sidesteps this inheritance by suspending the category label “social media” altogether. Crucially, the fictional frame does more than remove constraints; it makes certain desires and boundaries socially permissible to articulate. Wishes that might feel naive, awkward, or excessively earnest under a realistic prompt—wanting connection to feel like being in the same room, wanting visibility to be reversible, wanting strangers to remain strangers unless actively invited—become speakable when the task is to describe a magical world rather than repair an existing platform. In place of defensiveness, participants reason from felt social experience: what presence should feel like, what kinds of disclosure should be low-stakes, what boundaries should be socially intelligible.
A practical consequence is that FI generates common language between youth and adults. Much of the current tension in public discourse around social media stems from a mismatch in vocabulary: adults speak in terms of harm, addiction, and regulation, while youth often lack the conceptual tools to articulate what they want connection to feel like. FI bridges this gap by making youth priorities legible in terms that designers, researchers, and policymakers can act upon. When participants describe wanting “rooms” instead of “feeds,” or “neighborhoods” instead of “audiences,” they are not simply expressing preference for a feature; they are naming a different interaction grammar. That grammar, once articulated, becomes available for translation into design specifications, policy frameworks, and educational curricula. The contribution is not a single new feature, but evidence that a different organizing logic is possible and legible to users when the design frame allows it.
Finally, FI cultivates something often missing from conversations about social media: hope. The shift we observed from “I do not know what else there could be” to expansive, imaginative visions is not incidental. It reflects what happens when participants are given back authorship over the future, even temporarily. Past platforms were designed by adults with particular business models and assumptions baked in; youth imagining social worlds from scratch surfaces different priorities. The deeper problem is not that no one has tried to improve social media, but that most attempts operate within the same paradigm: adjusting privacy settings, adding screen time limits, tweaking algorithmic recommendations. These are repairs to a house whose floor plan is assumed fixed. FI invites participants to redraw the floor plan itself, supporting not only imagination but critical capacity: when youth can name what they want connection to feel like, and can describe the rules that would make that possible, they are better positioned to evaluate platforms critically without being trapped in cynicism. FI supports youth voice not only by eliciting ideas, but by cultivating the sense that different interaction futures are thinkable and therefore arguable.
5.4. Limitations and Future Work
This study is exploratory, and the design orientations we synthesize should be understood as generative rather than generalizable. Our participant demographics skewed toward women, and future work should examine how robust these orientations are across broader samples and cultural contexts, including youth with different relationships to gaming, spatial interfaces, and platform norms. Also, FI intentionally suspends feasibility and market constraints, which is a strength for defamiliarization but a limitation for direct translation into deployable systems. Some of the most compelling interaction logics we surfaced (spatially legible boundaries, ambient co-presence, rich memory-sharing) raise practical questions about accessibility, moderation, harassment, surveillance, and inequities in device access. Future work should prototype selectively, using low-fidelity and mid-fidelity implementations to evaluate whether the same felt qualities can be achieved under real constraints, and to identify new risks that emerge when imagined systems become operational.
@@ 6. Conclusion
This paper demonstrates that when youth are invited to imagine social connection outside the conceptual frame of “social media,” they articulate desires and interaction logics that diverge sharply from dominant platform paradigms. Through FI workshops with 23 participants aged 15–24, we surfaced six themes—intuitive social navigation, shared experiences, communal ambience, nuanced self-presentation, intentional engagement, and playfulness—that collectively describe what meaningful remote connection might feel like if it were organized around presence, place, and relationships rather than feeds, metrics, and passive consumption.
The findings carry implications for both research and practice. For researchers, this study offers evidence that design fixation in social media is partly a methodological problem: the term “social media” itself activates assumptions that constrain imagination. FI provides one way to interrupt those assumptions, enabling participants to reason from felt social experience rather than inherited platform templates. For designers and policymakers, the themes we identify suggest that the relational shortcomings of mainstream platforms are not inevitable features of mediated connection but outcomes of specific design choices—choices that could be made differently.
Appendix A Fictional Inquiry Materials
A.1. Miro Board for Fictional Inquiry Warm-Up
Figure 2 shows the Miro board template used during Fictional Inquiry sessions. Participants sorted their friends and family into Potter-inspired categories (houses, teams, old friends) to encourage thinking about relationships within the fictional context and enhance immersion in the design space. This activity preceded the main design exploration and helped participants transition from discussing their current social media experiences to imagining alternatives within the Hogwarts setting.
Appendix B FI Workshop Protocol